OpenStudy (anonymous):

INPUT GREATLY APPRECIATED! Without graphing the following equations, determine which one has the greater slope. Also, explain the process by which you came to your conclusion: a. 2y=5x-3 b. 5y=11x+3

OpenStudy (anonymous):

you need to put both equation in the form y = mx+b. since you have a coefficient on the y, simply divide both sides by it, and you'll get "m"

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

alright lets see what we hve

OpenStudy (anonymous):

I'm not really trying to solve I'm just figuring out which one would have the greater slope w/o graphing or anything.

OpenStudy (anonymous):

oohhhh ok hold on okay

OpenStudy (anonymous):

alright for A=5/2

OpenStudy (anonymous):

and for B=11/5

OpenStudy (anonymous):

those are the two slopes there

OpenStudy (anonymous):

Thanks! So basically you determined by solving?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

So can you please explain how exactly you solved?

OpenStudy (anonymous):

hehehe its a secret

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

use this website and i promise you youll get the right answer every time

OpenStudy (anonymous):

lol Thanks aminah!

OpenStudy (anonymous):

kk oh i have one question for you

OpenStudy (anonymous):

Ask away

OpenStudy (anonymous):

is that ur real pic

OpenStudy (anonymous):

Yes. Yes, it is.

OpenStudy (anonymous):

hmmmm hahahha ok :)

OpenStudy (anonymous):

? Why the laugh

OpenStudy (anonymous):

dont take this in a flirtatious sort of way but you are really cute

OpenStudy (anonymous):

Haha why thank you.

OpenStudy (anonymous):

hahah np ;)